OpenStudy (anonymous):

solve the initial value problem du/dt = (t+1)/sqrt(t); u(1) = 4

OpenStudy (amistre64):

\[\int\limits_{} \frac{t+1}{\sqrt{t}} dt\]

OpenStudy (amistre64):

split the fraction in two and work each side

OpenStudy (amistre64):

{S} t.(t^-1/2) + t^(-1/2) dt {S} t^(1/2) + t^(-1/2) dt

OpenStudy (amistre64):

t^(3/2) t^(1/2) ----- + ------ + C 3/2 1/2

OpenStudy (amistre64):

when t = 1; this should equal 4 according to the initial condition

OpenStudy (amistre64):

2/3 + 2 +C = 4 8/3 +C = 4 C = 4 - 8/3 = -16/3 if i did it right

OpenStudy (amistre64):

\[\frac{2t^2}{3}+\frac{2\sqrt{t}}{1}-\frac{16}{3}\]

OpenStudy (anonymous):

ahh thanks a ton. i was WAY overthinking it. used to other profs who constantly used all kinds of tricks and stuff so i was looking for something way more complicated. thanks again.

OpenStudy (anonymous):

though solving for c = 4/3, but otherwise it's golden

OpenStudy (amistre64):

:) youre welcome

OpenStudy (amistre64):

yeah, the integrating i can do.... addition? nah lol

OpenStudy (anonymous):

haha it's cool. i'm usually the same way- just apparently brain dead tonight